# Case Law Grounding: Using Precedents to Align Decision-Making for Humans and AI

### v3

@@ 4.2. RQ1: Comparing Case Law and Constitutional Grounding
Overall, these results suggest that for the vast majority of the tested groups and tasks, case law grounding did result in decisions that had statistically significantly higher accuracy compared to the ground truth. We also find AmItheAsshole in mod to be an outlier—while average accuracy was still higher the difference was not significant, which indicates that there may be communities and tasks where cases are not as effective. Coupled with results from RQ2b, we discuss in subsection 5.1 how limitations around the case law grounding process, such as challenges in situations where individual cases are longer and more complex, may have contribute to the reduced effectiveness of using cases as the main grounding mechanism for groups like AmItheAsshole.
@@ 4.3.1. RQ2a: Alignment Accuracy and Retrieval Window Size
In Figure 6 and Figure 7, we plot the simulated accuracy of CLG (case) in the human-led and LLM-prompting based processes at smaller retrieval window sizes than we used for RQ1 to evaluate whether smaller retrieval windows may be sufficient. In addition to plotting the scaling of case at different values of $k$ and presenting the rule condition for comparison at $k=15$, we also plot how 2 additional reference lines behave at different values of $k$: oracle and knn. As CLG uses binding precedents, the oracle reference line indicates an upper bound on accuracy given a certain retrieval window. For any case being judged, if the retrieval window contains any past case with the ground truth decision, oracle will always produce the correct answer, otherwise it will always be incorrect. We can observe that at small retrieval windows, it is often the case that across the entire set of judged cases, even the oracle cannot achieve perfect accuracy, as for some cases no correct selection of precedents exists. However, as $k$ becomes larger, we see oracle more consistently achieving perfect accuracy. The knn reference line represents a strategy that always selects all precedents and relies on the tie-breaking mechanism in CLG to produce the final decision. Deviations between case and knn indicates the effect of contributions from the decision-maker as a result of precedent selection.
@@
With both human and LLM processes, we did not find any consistent correlation between $k$ values and accuracy across the tested groups and domains, nor did we find statistically significant differences between accuracy of neighboring $k$ values in the set of values tested. This suggests that for many of the individual groups and domains tested, smaller $k$ values may be sufficient to achieve a similar level of accuracy. However, looking across groups we also observed the following behavior patterns that we believe suggest against picking very small $k$ values to reduce decision cost, especially since it is not possible to evaluate scaling in practice when ground truth is not available:
@@
Performance plateau behavior. In most of the groups and domains (excluding the outlier AmItheAsshole we observed in RQ1 evaluations), accuracy tends to eventually reach a plateau where further scaling of $k$ contributes little to accuracy. However, the speed at which this is reached varies by the group and domain: For example, groups like seen=Y,targeted=N or Coronavirus seem to reach a plateau at $k=3$ and $k=5$, while Games still sees accuracy improvements at $k=10$. While this behavior not too surprising by itself—with more dissimilar cases retrieved, we should expect additional cases to be less and less relevant to the final decision—there does not seem to be a consistent $k$ for reaching the plateau across groups and tasks. We hypothesize that this variation across groups is likely to do with aspects like the “density” (average distance) of past cases that can be retrieved and distributional properties of past cases around the judged cases.
@@
Unreliable for retrieval windows that are too small. An additional observation is that, at very small values of $k$ such as $k=1$ or $k=3$, performance can be relatively unpredictable. In some groups, like seen=Y,targeted=Y and Coronavirus, we see initial drops in performance at lower $k$ values. While this drop is not significant when viewed across all the judged cases, it does indicate that accuracy is much more sensitive to $k$ when $k$ is small—moving between retrieval thresholds can impact accuracy on some portions of cases. We hypothesize that this may be due to case-based approaches naturally being more sensitive when the retrieval window is low, as small changes could result in relevant precedents not making it past the threshold or immediate local neighborhoods primarily comprising of similar but less relevant cases.
We discuss the implications of some of these observations in subsection 5.1, noting potential areas of future work that may provide signals for optimal $k$ selection without relying on an annotation-intensive process like how we are evaluating RQ2a, where an examination of scaling is conducted using a subset of cases where ground truth is known. Finally, we also found that the AmItheAsshole group (which was an outlier before) still to exhibits distinctive behavior—we see accuracy first improve as the window size increases but then fall back down (though still not to a statistically significant degree, with even the top performing $k$ still not being statistically significantly higher performance than rule, $p=0.079>0.05$). However, as this behavior happens at small $k$ ($k=3$ and $5$), we believe this is reflects the earlier observation of unpredictable performance when $k$ is small.
@@ 4.3.2. RQ2b: Binding Precedents versus Cases as Examples
To examine this, during the human annotation process, we not only collected participants’ judgments on whether each precedent applied, we also asked participants to give a final decision on the judged case after reviewing the list of retrieved prior cases without being constrained to the consensus of the the cases they selected as precedents (as noted in subsection 3.3). In Figure 4, we include an additional condition—example—to reflect the accuracy of this configuration. Comparing the accuracy produced by non-binding precedents (example) to that of constitutional grounding (rule), we note the absence of statistically significant differences (paired t-test, $p=0.635,0.360,0.458,0.266>0.05$ for seen=Y,targeted=N, seen=Y,targeted=Y, Coronavirus, Games) in most of the groups, despite the fact that we did observe statistically significant improvements to accuracy when using binding precedents (case). Here, again, AmItheAsshole is an outlier, with example showing a statistically significant improvement over rule (+20.0 p.p, $p=0.020<0.05$) even though binding precedents (case) did not show significant improvements.
4.4. RQ3: Assessing the LLM Workflow with Larger Datasets
(a) LLM accuracy on the additional cases in tox
(b) LLM accuracy on the additional cases in mod
Figure 8. LLM performance and scaling on additional cases for the tox and mod domains. Reference lines for oracle and knn are produced in the same way as Figure 6. rule also reflects LLM performance using constitutional grounding on these cases. Retrieval still done at $k=15$ with lower $k$ windows simulated post-hoc.
[…]
Line plots
Finally, in our last evaluation, we use a larger set of cases (that we did not evaluate the human-led process on) to evaluate the LLM-prompting based process for CLG, examining whether our observations from RQ1 and RQ2 still hold (Figure 8). At the same $k=15$ retrieval window, we find that across all 6 groups at this larger scale, case law grounding (case) consistently outperformed constitutional grounding (rule) when using the prompting version: For the tox domain, we observed +20.7 p.p. for seen=Y,targeted=N (paired t-test, $p=0.0002<0.05$), +24.8 p.p. for seen=Y,targeted=N (paired t-test, $p<10^{-4}<0.05$), and +13.6 p.p. for seen=N,targeted=N (paired t-test, $p=0.016<0.05$). For the mod domain, we observed +9.2 p.p. for Coronavirus (paired t-test, $p=0.0418<0.05$), +13.3 p.p. for AmItheAsshole (paired t-test, $p=0.0047<0.05$), and +30.0 p.p. for Games (paired t-test, $p<10^{-5}<0.05$). We also observe similar trends (plateaus and more variations at low $k$) with respect to scaling over $k$ as we did for RQ2a, with the only major difference being that with this larger dataset, we do now see improvements sufficient for statistical significance even with the AmItheAsshole that was previously an outlier without significant results.
@@ 5. Discussion
In this section, we will discuss some of the limitations of case law grounding and how they may have contributed to some of our observations in our experiment results. Then, we will discuss implications and considerations around the human-led decision process versus LLM prompting. Finally, we will discuss how CLG and constitutional grounding can be complementary solutions to the problem of grounding subjective decisions to group or community norms and preferences, and how we can view the two processes within the framework of policy-making in communities.
5.1. Limitations of Cases as Grounding Mechanisms
As we found in our experiments, while CLG was effective compared to constitutional grounding in most of the group decisions evaluated, there were also situations where CLG did not perform significantly better than constitutional grounding. One possible explanation for this may lie in limitations of cases themselves when it comes to grounding decisions. While the concrete nature of cases means it may be less ambiguous for decision-makers (human or LLM) to assess each case, the specificity encoded in cases can also have several drawbacks.
For one, cases contain details and context that support making the decision, meaning they can be longer and more complex. Since CLG uses similarity-based metrics for retrieval, longer cases means higher chances of spurious similarities being picked up by the retrieval process, degrading the quality of set of retrieved cases. Compared to other groups in mod, AmItheAsshole is a community around interpersonal conflicts, and as a result some comments contain much more descriptive details while others do not. Additionally, because these comments contain complex context, it is likely that even when a precedent is relevant, it may not be associated with the right “keep”/“remove” decision due to other details of the case. Considering the observation that non-binding precedents performed significantly better while binding ones did not, this limitation may have been an important factor.
Furthermore, CLG performance can depend on the (local) density of past decisions around each new decision in the retrieval window. As we observe in RQ2a, CLG can perform well for some groups even with small retrieval windows, thus saving on costs (in human effort or model computation). However, window sizes that are too small risk missing relevant precedents while too big of a retrieval window risks introducing irrelevant and confusing prior decisions. While our evaluation setup in RQ2a conveniently allows us to evaluate different values of $k$, in a practical deployment we would not have access to the ground truth decisions and thus may end up with a sub-optimal $k$.
Finally, beyond mechanistic limitations of cases, using case law as the main grounding mechanism can also incur social or emotional burdens on human decision-makers. In the domain of the content moderation, prior work (Steiger et al., 2021) has noted that simply assessing harmful and disturbing content imparts an emotional toll on moderators. With case law grounding approaches, we may end up exposing human decision-makers to a higher amount of harmful content during their precedent selection compared to constitutional grounding. While we took measures to remove identifiable information in our experiment, some annotators still expressed concerns around privacy when showing full text of past posts or comments.
We believe there are many opportunities for future work to address these limitations, ranging from using new metrics to estimate optimal $k$ to LLM-based tools rewrite cases while balancing removal of irrelevant or identifiable context.
5.2. Large Language Models for Community Decision-Making
In our experiments, we observed that despite not being specifically prompted or tuned for a particular group or community, the LLM version of CLG (case) still performed relatively well, in certain cases slightly outperforming human annotators when it comes to average accuracy (though overall we did not observe statistically significant different accuracy between CLG with human decision-makers or CLG with LLM prompting, with both performing on-par with each other). Looking at the groups that saw the most difference between human and LLM performance, we note a couple of potentially interesting differences:
In AmItheAsshole, the LLM process produced slightly higher accuracy for the CLG (case) condition. For many of the cases where the LLM produced a correct decision while the human decision-maker did not, the set of retrieved past cases often contained more complex and longer cases, coupled with the fact that a $k=15$ retrieval window was used during annotation, this possibly suggests that for CLG with more complex cases, human decision-makers may be making more-than-expected errors due to possible fatigue. In Games, we also observe that LLMs did much worse with constitutional grounding (rule) than human decision makers. Examining the constitutional guidelines, we found that, unlike other groups, Games had a very long and complex set of community guidelines (spanning a total of 27 individual rules with detailed descriptions for each as opposed to around 10 rules for Coronavirus and AmItheAsshole). We hypothesize that in this case, human decision-makers were better able to intuit many of the complex rules quickly, whereas the much longer and detailed set of instructions may have made it harder for LLMs to apply the correct criteria.
Overall, while we did not find any clear patterns that suggest LLM prompting makes more specific types of failures (at a group level) compared to human decision-makers when using CLG, we do notice that human decision-makers did slightly better when using constitutional grounding than LLMs. We hypothesize that this may be due to a combination of the fact that human participants were drawn from members of the relevant groups / communities, and thus may have internalized some of the norms and preferences of the group, coupled with more familiarity when consuming community guidelines which were written with human understanding in mind. Given these observations, we believe that if groups and communities wish to start adopting LLMs to automate decision-making, following more of a case law approach rather than supplying existing constitutions made for humans to a LLM is likely to result in better results. For the groups and communities we evaluated, past decisions seem to be a much stronger signal for grounding future decisions than their constitutional guidelines, though this may not generalize to more marginalized groups or communities around more sensitive topics.
5.3. The Complementarity of Constitutions and Case Law
In this work, we explored how a case law inspired approach to grounding the decision-making process in socially defined contexts can lead to closer alignment towards group and community preferences when compared to a purely constitutional approach using existing guidelines. However, constitutions and case law do not oppose each other in practical applications, and in legal systems it is not uncommon to have mixed systems that utilize both in conjunction (Palmer, 2012). Thus when looking at decision-making within socio-techincal systems, it is also useful to see both constitutions and case law as complementary means for grounding decision-making, with their own strengths and weaknesses.
Take the example of content moderation in online communities: It’s common to see communities organically utilize constitutional grounding in the form of “community guidelines”. However, others have also observed that community guidelines are not always enforced consistently by moderators in the wild, and not all community members agree with all the specific norms and criteria described by the community guidelines (Jhaver et al., 2019a). Nonetheless, community guidelines are still a helpful means for communities to convey values and norms to both existing members and newcomers alike, even if they may not have the comprehensiveness required for making certain situational decisions. Here is where we believe case law and constitutions can act as complements to each other: The concise abstractions offered by constitutions help ground general values and preferences when making decisions, while the specificity and nuance offered by precedents help clarify how to consistently apply these values in-context. In fact, during our collection of community guidelines on Reddit for our study, we also observed communities including community-specific examples within their constitutional guidelines as a tool for clarifying concepts or making situational exceptions—both aspects we believe a case-based approach like CLG excel at.
Having constitutions and case law as complementary components can also be useful in situations that demand socially-appropriate decisions without complete human supervision. The recent commercial deployments of large language models (LLMs) as general purpose AI agents has also brought along high-profile cases of models failing to consider the social appropriateness of responses (Weidinger et al., 2021). Unlike with content moderation, it is impractical to have human decision-makers determine appropriateness for each user query and response as they come. Accordingly, providers of these AI agents have been working on ways to ensure that the outputs of such systems also follow the norms and expectations of broader society based on constitutional approaches as well. One such example of this is with Constitutional AI (Bai et al., 2022), where the general public is tasked with helping define universal values into a constitution that is then incorporated into model training and tuning. Here again, the case-based approach in CLG could provide additional utility where the LLM-backed decision workflow would allow both constitutions and retrieved in-context examples to be applied during inference time to determine appropriate responses or strategies for responding, in a way similar to recent work with retrieval augmentation for generative models and in-context learning (Lewis et al., 2020b; Lin et al., 2024). This could help bridge potential issues where models may “interpret” human-produced guidelines differently, by also introducing relevant human-made decisions.
5.4. Addressing Consensus and Evolution of Preferences
In our current implementation of CLG, we use a simple approval voting mechanism to resolve conflicts around the final decision given the precedent set. However, in many situations conflicts can be a signal that past cases may be insufficient to ground the current decision and that we must clarify the decision bounds. While CLG currently does not include a mechanism for this kind of consensus-building, we can look towards prior work for possible implementations that could be incorporated into CLG. For example, when conflicts demand clarity on new decision criteria, voting systems like digital juries (Pan et al., 2022) could be used to engage opinions of different stakeholders. On the other hand, when conflicts arise from disagreements when reasoning about how to apply past cases, deliberation may be the tool to resolve them (Chen et al., 2019).
Another common criticism of case law in general centers around issues with the doctrine of stare decisis (Douglas, 1949; Kozel, 2010). Since societal and community values tend to evolve over time (Lin et al., 2017), if future decisions must always be bound to past decisions, then case law decisions would grow increasingly detached from current values. To address this in the long term, case law grounding also needs ways to retire or overturn past precedents. Some have even pointed to this flexibility as a benefit of case law when it comes to quickly adapting to shifting preferences of the public (Cardozo, 2011).
With relatively minor modifications, the CLG process could be adapted to support decision-making over a long period of time by automatically retiring old cases and decisions that have become “stale” and rarely cited and overturning past decisions through a simple escalation process during precedent selection—in addition to indicating relevance, decision-makers could also flag precedents whose decisions they disagree with. By additionally including book-keeping around precedent dependency chains, retired or overturned cases can also automatically trigger re-evaluations of past decisions that depended on them.
5.5. Policy-Making through Case Deliberation
One aspect that differentiates of constitutional grounding from case law is how constitutions often facilitate policy-making (Zhang et al., 2020) in addition to being a way to ground decisions on policy. Groups and communities often use constitutions and guidelines as a way to formalize their norms and preferences, with the process of making constitutions a potentially important component to reflect on the values of the group or community (Butler et al., 2008). In these respects, constitutions can have the advantage of being concise, and thus easy to form opinions on, as well as abstract, allowing coverage of a large part of the decision space with relatively low effort. For larger communities, like that of Wikipedia, there can be complex documented procedures that can be triggered by members to evaluate proposed policy against current practice and used to formalize core policy values into concise, applicable rules or even implemented as bots and algorithms (Hwang and Shaw, 2022). Indeed, the misalignment observed between decisions and guidelines could in part be attributed to the sometimes aspirational nature of constitutions to reflect ideals.
In contrast, while case law can be easy to bootstrap initially—in many cases like content moderation, past decisions are readily available to be used as precedents—building a precedent-based policy solely from naturally encountered decisions can be inefficient, as each precedent only addresses a small part of the decision space. Precedents themselves also offer limited insight into higher-level patterns like community or group values, as each decisions on cases are specific and contextual in nature. However, cases can still serve as a useful medium for policy-making in a different way, by helping groups and communities better consider nuances in decision bounds and “edge cases”. When policy-making only focuses on high-level constitutions, decision-makers are left to interpret ambiguities themselves when policies need to be applied. Because cases are specific in nature, if we can identify or even synthesize cases that lie within the ambiguous spaces between high-level constitutional guidelines, we could surface them during community policy-making processes, where they could contribute to more clarity for the resulting constitution, or, when combined with democratic processes, become binding precedents for controversial situations. We envision complete future applications of case law grounding in community and group settings to also involve case-based policy-making components. Specifically, a case-based policy-making process could analyze existing decisions and precedents to identify areas where policy ambiguity is high or where disagreements are common. These results could serve as signals for selecting or synthesizing challenging cases to collect community inputs on, with the end goal of producing both precedents that directly define decision bounds and potential intermediate probes for building and refining constitutional guidelines.

### v4

@@ 4.1.1. Tasks and Datasets
4.1.2. Communities and Groups
A major challenge in making decisions over subjective and socially-defined concepts lies in the locality of preferences—criteria for the same decision task like toxicity rating or moderation can differ considerably between groups and communities (Koshy et al., 2023). To capture this, we further split each task domain into specific task instances defined by the group or community the decisions should align to.
For the mod domain, we selected 3 large communities from Reddit that were captured by the NormVio dataset. For the tox domain, demographic backgrounds of raters can contribute to significant preference differences, so we take the most significant two (prior exposure and observers versus victims) such demographic dimensions identified by prior work (Kumar et al., 2021). Overall, this resulted in 5 group-defined task instances, details of which we document in subsection A.2.
4.1.3. Participant Recruitment and Crowdsourced Annotation
We recruited participants for the mod tasks directly from the respective subreddits. For tox, we recruited participants on Prolific, a crowdsourced survey platform. In both domains, we used a screening survey to identify participants that would be a good fit. Human participants in both conditions were paid a base rate of \$15 for completing our annotation study (30-45 minutes), with Prolific participants being paid through the platform and Redditors being paid in the form of an Amazon Gift Card. In total, we were able to recruit $n=60$ participants from Prolific and $n=35$ participants from Reddit. We document the additional details around recruitment in subsection A.3.
4.1.4. Additional Setup Details
To facilitate evaluations such as the simulation study or examination of binding-ness of precedents, we made additional configurations to the annotation task setup. These details, as well as general configuration parameters, are documented in subsection A.4.
4.2. RQ1: Comparing Case Law and Constitutional Grounding
@@
Overall, these results suggest that for the vast majority of the tested groups and tasks, CLG resulted in decisions that had statistically significantly higher accuracy compared to constitutional grounding. We find AmItheAsshole in mod to be an outlier—while average accuracy was higher, the difference was not significant, indicating that grounding with cases can be less effective in some situations.
@@ 4.3.1. RQ2a: Alignment Accuracy and Retrieval Window Size
In Figure 6 and Figure 7, we plot the simulated accuracy (subsection A.4) of CLG (case) in the human-led and LLM-prompting based processes at smaller retrieval window sizes than we used for RQ1 to evaluate whether smaller retrieval windows may be sufficient. In addition to plotting the scaling of case at different values of $k$ and presenting the rule condition for comparison at $k=15$, we also plot how 2 additional reference lines behave at different values of $k$: oracle and knn. The oracle reference line indicates an upper bound on accuracy given a certain retrieval window. For any case being judged, if the retrieval window contains a past case matching ground truth, the oracle will have perfect accuracy. We observe that at small retrieval windows, it is often the case that across the entire set of judged cases, even the oracle cannot achieve perfect accuracy, as for some cases no correct selection of precedents exists. However, as $k$ becomes larger, we see oracle more consistently achieving perfect accuracy. The knn reference line represents a strategy that always selects all precedents and relies on the tie-breaking mechanism in CLG to produce the final decision.
@@
In both human and LLM processes, we did not find any consistent correlation between $k$ values and accuracy across the tested groups and domains, nor did we find statistically significant differences between neighboring $k$ values when it came to final accuracy, suggesting that for many of the individual groups and domains tested, smaller $k$ values may be sufficient to achieve a similar level of accuracy.
However, looking across groups we also observed the following behavior patterns that we believe suggest against picking very small $k$ values to reduce decision cost, especially since it is not possible to evaluate scaling in practice when ground truth is not available:
@@
Performance plateau behavior. In most of the groups and domains (excluding the outlier AmItheAsshole from before), accuracy tends to reach a plateau where further scaling of $k$ contributes little to accuracy. However, the speed at plateauing varies by the group and domain: For example, groups like seen=Y,targeted=N or Coronavirus seem to reach a plateau at $k=3$ and $k=5$, while Games still sees accuracy improvements at $k=10$. While this behavior is not too surprising by itself—with more dissimilar cases retrieved, we should expect additional cases to be less and less relevant to the final decision—there does not seem to be a consistent $k$ that ensures plateauing across groups and tasks. We hypothesize that this variation across groups is likely to do with aspects like the “density” (average distance) of past cases that can be retrieved and distributional properties of past cases around the judged cases.
@@
Unreliable for retrieval windows that are too small. We also observe that very small values of $k$ such as $k=1$ or $k=3$ result in unpredictable performance. For seen=Y,targeted=Y and Coronavirus, we even saw drops in performance at lower $k$ values. While these drops are not significant when viewed across all judged cases, they indicate that accuracy is more sensitive to $k$ when $k$ is small—moving between retrieval thresholds can impact accuracy on some portions of cases. We hypothesize that this may be due to case-based approaches naturally being more sensitive when the retrieval window is low, as small changes could result in relevant precedents not making it past the threshold or immediate local neighborhoods primarily comprising of similar but less relevant cases.
Ultimately, we don’t find generalizable correlations between changes in the retrieval window $k$ and final performance, only noting that for any task / group, $k$ should be sufficiently large so as to reach the performance plateau; otherwise (when $k$ is too small) performance can suffer.
@@ 4.3.2. RQ2b: Binding Precedents versus Cases as Examples
To examine this, during the human annotation process, we not only collected participants’ judgments on whether each precedent applied, we also asked participants to give a final decision on the judged case after reviewing the list of retrieved prior cases without being constrained to the consensus of the cases they selected as precedents (as noted in subsection 3.4). In Figure 4, we include an additional condition—example—to reflect the accuracy of this configuration. Comparing the accuracy produced by non-binding precedents (example) to that of constitutional grounding (rule), we note the absence of statistically significant differences (paired t-test, $p=0.635,0.360,0.458,0.266>0.05$ for seen=Y,targeted=N, seen=Y,targeted=Y, Coronavirus, Games) in most of the groups, despite the fact that we did observe statistically significant improvements to accuracy when using binding precedents (case). Here, again, AmItheAsshole is an outlier, with example showing a statistically significant improvement over rule (+20.0 p.p, $p=0.020<0.05$) even though binding precedents (case) did not show significant improvements.
@@ 5. Discussion
5.1. Limitations of Our Implementation
In this work, we explored a relatively simple implementation of case law for grounding decisions where reasoning around cases is abstracted simply into a selection process. However, in the legal space, there are many specific conventions around how precedents are actually used (Ashley, 1991; Sunstein, 2021). While the degree to which case-based and analogical reasoning can be applied is still openly debated in the legal realm (Caputo, 2024), there are opportunities to improve the value of grounding cases by making use of full-fledged analogical reasoning.
Furthermore, we also observed differences between groups / communities in how CLG affected decision alignment accuracy in our experiments, with notable cases like AmItheAsshole seeing a very different performance pattern. While we don’t have sufficient results to examine this more comprehensively, we do hypothesize that the permissiveness and concreteness of the community’s internalized criteria itself may be at play. Compared to other communities examined, AITA was much more permissive about what content is acceptable: even in ambiguous cases (like suspected fabricated stories), enforcement of moderation tended to lean lenient. This inherent acceptance of ambiguity could contribute to less visible differences between constitutional and case law style grounding, which specifically focuses on challenges with ambiguous concepts.
5.2. Integrating CLG into Decision-Making Ecosystems
While we primarily compared CLG and constitutional grounding in this work, both approaches offer unique values in the broader ecosystem of decision-making. When we look at prior examinations of community governance (Hwang and Shaw, 2022), we often see a combination of rules and cases being used throughout the process—with cases contributing to “legislative” processes that produce the very rules and guidelines that are adopted while also being applied in debates over ambiguous cases. Similar patterns of adopting both abstractions (constitutional rules) and specifics (case-based analogical reasoning) are common in high-stakes decision-making tasks such as with jurisprudence (Caputo, 2024).
Ultimately, constitutional and case-law grounding are suited to different challenges in decision-making, and with this work, our goal is to bring to light the need for more case-based approaches around the challenging decision-making tasks like those involved in online community governance or in areas like AI alignment. We believe there is much to be explored around future processes that combine both forms of grounding to efficiently produce high-quality decisions—both in human-led and AI automated settings.